OpenStudy (hba):

If (2x-1) +(3y+2)i=5-4i,then(x,y) is

OpenStudy (hba):

@hartnn

hartnn (hartnn):

just equate real and imaginary parts 2x-1=5 3y+2=-4

hartnn (hartnn):

now thats pretty simple to solve

OpenStudy (hba):

(3,-2) @hartnn

hartnn (hartnn):

yup

OpenStudy (hba):

thanks :)

OpenStudy (hba):

@hartnn 3x+(y-2)i=(5-2x)+(3y-8)i

OpenStudy (hba):

I got (1,3)=(x,y)

hartnn (hartnn):

yes, correct

OpenStudy (hba):

Yuppie :)

hartnn (hartnn):

now you are good at complex numbers :)

OpenStudy (hba):

Maybe :D